OpenStudy (anonymous):

(8,0), m=-3 Find an equation of the line that passes through the given point and has the given slope. Express your answer in slope intercept form

OpenStudy (aroub):

y=mx+b, m=-3, (8,0) 0=-3(8)+b 0=-24+b 24=b y=-3x+24

OpenStudy (anonymous):

first put it in point slope form: 0-y=-3(8-x) distribute: 0-y=-24+3x rearrange and simplify: -y=3x-24 y=-3x+24